# From Image to UML: First Results of Image-Based UML Diagram Generation using LLMs

### v1

@@ 5.1.2 Threats to external validity
The nondeterministic nature of LLMs might affect the validity of the interpretation of the results. This is further the case due to the small sample number of attempts used to evaluate the LLMs per prompt and per image. Although, ignoring the results with syntax errors, the number of errors in the generated PlantUML code per LLM seemed to vary very little, showing consistent performance for each example.
Generalization to other LLMs
@@
Although the results of the experiment reflect the performance of the used visual LLMs, the results do not generalize to other visual LLMs. Indeed, the experiment itself has shown that some LLMs are more adequate for the explored task than others. This means, rather than saying that all visual LLMs are capable or not capable of transforming images of UML diagrams to a computer readable format, it always depends on the LLM. Thus, if another model is to be used, tests are required.
@@ 6 Tool support
In a first step towards low-modeling, the BESSER [^6] https://github.com/BESSER-PEARL/BESSER platform includes a UML class diagram image to PlantUML converter. BESSER, which focuses on the efficient development of smart software, implements an LLM interface that currently supports GPT-4V via the OpenAI API. Beyond the transformation to PlantUML, the library also offers a transformation to the BESSER Universal Modeling Language (B-UML), that is the UML-inspired language of the BESSER low-code platform. It aims to leverage the advantages of UML while also having the freedom to integrate and extend the language with other (meta)models based on the requirements. The transformation to B-UML enables an immediate usage of the BESSER generations, effectively completing the pipeline from image, to computer readable format, and to software. The BESSER-examples [^7] https://github.com/BESSER-PEARL/BESSER-examples repository contains guidelines on how to use the image to UML functionality.
@@ 7 Conclusion and future work
In the future, we plan to conduct qualitative studies, inviting human participants, software engineering experts or computer science students, to evaluate the usability and usefulness of the concept and results. Additionally, we would like to evaluate the ability of LLMs to transform other types of UML diagrams, such as state diagrams or use case diagrams, but also, more generally, non-UML diagrams often used in software engineering such as entity relationship diagrams.
On the technical side, we want to improve the process by automatically checking some correctness properties of the recognized models as part of a pipeline that would automatically trigger new calls to the LLMs asking them to revise the detected mistakes. A second LLM could also be used for this purpose. This LLM-as-judge would evaluate the results and decide whether a new generation is needed.
Finally, we would like to explore the usefulness of this transformation in an educational context where the LLM could play different roles. For instance, one could instruct the LLM to act as a mentor or teacher to the students using it to solve problems. Based on the received input from the students, the LLM could either immediately correct the given image and give feedback or even return the generated UML class diagram alongside with feedback on aspects it perceives as wrong and nudge the students to correct them, acting as an educational agent.
Acknowledgments
This project is supported by the Luxembourg National Research Fund (FNR) PEARL program, grant agreement 16544475

### v2

@@ 5.2 Threats to external validity
Nondeterministic nature of LLMs and sample size. The nondeterministic nature of LLMs might affect the validity of the interpretation of the results. This is further the case due to the small sample number of attempts used to evaluate the LLMs per prompt and per image. Although, ignoring the results with syntax errors, the number of errors in the generated PlantUML code per LLM seemed to vary very little, showing consistent performance for each example.
@@
Generalization to other LLMs. Although the results of the experiment reflect the performance of the used visual LLMs, the results do not generalize to other visual LLMs. Indeed, the experiment itself has shown that some LLMs are more adequate for the explored task than others. This means, rather than saying that all visual LLMs are capable or not capable of transforming images of UML diagrams to a computer readable format, it always depends on the LLM. Thus, if another model is to be used, tests are required.
@@ 6 Tool support
In a first step towards low-modeling, the BESSER platform includes a UML class diagram image to PlantUML converter. BESSER [17], which focuses on the efficient development of smart software, implements an LLM interface that currently supports GPT-4V via the OpenAI API. Beyond the transformation to PlantUML, the library also offers a transformation to the BESSER Universal Modeling Language (B-UML), that is the UML-inspired language of the BESSER low-code platform. It aims to leverage the advantages of UML while also having the freedom to integrate and extend the language with other (meta)models based on the requirements. The transformation to B-UML enables an immediate usage of the BESSER generations, effectively completing the pipeline from image, to computer readable format, and to software. The BESSER-examples [^6] https://github.com/BESSER-PEARL/BESSER-examples repository contains guidelines on how to use the image to UML functionality.
@@ 7 Conclusion and future work
In the future, we plan to conduct qualitative studies, inviting human participants, software engineering experts or computer science students, to evaluate the usability and usefulness of the concept and results. Additionally, we would like to evaluate the ability of LLMs to transform other types of UML diagrams, such as state diagrams or use case diagrams, but also, more generally, non-UML diagrams often used in software engineering such as entity relationship diagrams. On the technical side, we want to improve the process by automatically checking some correctness properties of the recognized models as part of a pipeline that would automatically trigger new calls to the LLMs asking them to revise the detected mistakes. A second LLM could also be used for this purpose. This LLM-as-judge would evaluate the results and decide whether a new generation is needed. Finally, we would like to explore the usefulness of this transformation in an educational context where the LLM could play different roles. For instance, one could instruct the LLM to act as a mentor or teacher to the students using it to solve problems. Based on the received input from the students, the LLM could either immediately correct the given image and give feedback or even return the generated UML class diagram alongside with feedback on aspects it perceives as wrong and nudge the students to correct them, acting as an educational agent.
Acknowledgements.
This project is supported by the Luxembourg National Research Fund (FNR) PEARL program, grant agreement 16544475 and the CLIMABOROUGH project, funded by the European Union under the grant agreement 101096464.